# Learning-assisted Stochastic Capacity Expansion Planning: A Bayesian Optimization Approach

### v3

@@
Abstract
@@
Index Terms:
planning, power systems, stochastic optimization, mixed integer linear programming, machine learning, uncertainty
I Introduction
I-A Motivation
@@
I-B Time Series Aggregation
@@
Figure 1: Effect of varying the number of representative days on the CEP objective evaluated over 14 out-of-sample supply-demand projections using a single projection solution for a joint power-gas CEP (Sec. IV).
I-C Contributions
We consider two-stage stochastic CEP for coordinated planning of multiple energy vectors. We develop a learning-assisted approach to compute investment decisions that yield the lowest combined investment and operational cost in expectation over multiple year-long projections of energy demand and VRE availability (Sec. II). Rather than solve an extensive form stochastic program, which is intractable due to the large-scale and combinatorial nature of the problem, our approach designs and solves a surrogate CEP instantiated using a smaller set of representative periods. Note, however, that there can be many potential surrogate CEPs, each yielding different planning outcomes. Our key contribution is to leverage a Bayesian optimization (BO) approach to efficiently search the space of RPC hyperparameters. We demonstrate that these automatically tuned hyperparameters help identify planning outcomes with the lowest expected total cost (investment plus operational) over out-of-sample supply-demand projections (Sec. III).
@@ I-C Contributions
Our BO-assisted approach can be viewed as a “learning to configure” approach [22] in that it learns to select hyperparameters specifying the time series aggregation heuristic for surrogate model construction. In this regard, our approach is distinct from recent works that use offline learning to improve optimization heuristics [23, 24] as it can be deployed without pre-training on a large number of problem instances (Fig. 2). Specifically, our approach quickly identifies low-cost investment decisions by learning to identify promising hyperparameter configurations for a single problem instance through an iterative process of instantiating surrogate problems, solving them, and evaluating the resulting investment decisions over a range of validation supply-demand projections. We demonstrate the effectiveness of this approach for joint power-gas generation and transmission expansion planning for New England with 20 yearly supply-demand projections (Sec. IV).
II Problem Formulation
II-A Capacity Expansion Problem
@@
We assume that (II-A) is feasible for any first-stage decisions, $x$, i.e., load shedding is unconstrained and incurred as a cost. Additionally, we note that each period, $t\in\mathcal{T}$, does not necessarily correspond to the planning resolution of energy systems. For instance, while hourly time resolution is usually used for electric power planning, daily operating resolution can be more suitable for gas system planning [3]. Following [13, 26, 27], we assume representative periods to denote daily time spans.
II-B Surrogate Capacity Expansion Problem
@@
II-C Representative Period Clustering Problem
The choice of representative days is crucial to achieving a computationally light surrogate model that well-approximates (II-A). In other words, the investment decisions $x$ resulting from $\text{CEP}_{\text{S}}(\mathcal{T}_{R},\mathcal{W})$ should approximately minimize (1a). Existing literature, including [29, 26, 14, 12], propose clustering-based approaches for identifying such representative periods. Here, we consider the following k-medoids clustering model, which we refer to as the representative period clustering (RPC) problem:
@@ II-C Representative Period Clustering Problem
where $\mathcal{C}\coloneqq(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ denotes the representative day clusters with corresponding medoids $(m_{1},\dots,m_{k})$, $D(m_{i},t)$ denotes the “distance” between medoid $m_{i}$ and period $t$, and $\mathcal{T}_{E}\subset\mathcal{T}_{R}$ denotes a set of “extreme days,” or days exhibiting particularly hard-to-satisfy demands, with corresponding weights $w_{i}=1$ for all $i\in\mathcal{T}_{E}$. One can apply an out-of-the-box clustering algorithm [30] to solve (II-C) and obtain the representative day set and weights respectively as $\mathcal{T}_{R}=(m_{1},\dots,m_{\theta_{R}})$ and $\mathcal{W}=(\frac{1}{|\mathcal{S}|}|\mathcal{C}_{1}|,\dots,\frac{1}{|\mathcal{S}|}|\mathcal{C}_{\theta_{R}}|)$. Here, we focus on the central issue of selecting hyperparameters for the RPC problem.
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We consider $\theta_{B}\in[0,1]$ to be a hyperparameter that weighs the relative importance of demand (contained in ${q}_{t}^{\omega}$) and supply-side features such as renewable capacity factors (contained in ${B}_{t}^{\omega}$). Importantly, different choices of $\theta_{B}$ impact RPC outcomes by influencing the daily supply- or demand- patterns captured in the representative day set. Selecting $\theta_{B}$ is not obvious. More generally, more than two groups of parameters may need to be weighed when multiple VRE technologies and energy vectors are considered, which is the case for our computational study (Sec. IV).
@@
Figure 4: Pearson correlations between supply-demand parameter groups for our computational study (Sec. IV). Each row/column of pixels corresponds to one hour of the day (averaged over system nodes) except for NG, where pixels correspond to daily nodal loads. For each pair of parameter groups, pixel $(i,j)$ shows the correlation of parameter $i$ in the first group with parameter $j$ in the second group.
@@
III Solution Approach
To address this challenge systematically and for a range of planning contexts, we propose searching over a continuous space of RPC hyperparameters using Bayesian optimization (BO), a derivative-free search strategy for black-box functions that has been applied with success in hyperparameter tuning for ML applications [32]. At a high level, our BO-assisted approach learns to minimize $f(\cdot)$, the function mapping RPC hyperparameters to CEP objective values, by performing a series of function evaluations. As is the case in many BO applications, $f(\cdot)$ does not admit an analytic expression and can only be evaluated through a series of complex operations (i.e., solving optimization problems). In each iteration of our approach, a function estimate – a Gaussian process (GP) regression model – is re-estimated using the most recent function evaluation, and new candidate hyperparameters are identified according to their estimated potential for yielding a low CEP objective and improving the GP model estimate. Fig. 6 illustrates our solution approach and shows how hyperparameters $\theta$ are evaluated (Sec. III-A) and how subsequent candidate hyperparameters are selected (Sec. III-B) in each iteration.
III-A Function Evaluation
@@
1) Clustering the sample set. We solve (II-C) using a sample set constructed from $\Omega^{\text{train}}$ to identify a set of representative days (i.e., cluster medoids) and their corresponding weights, $\mathcal{T}_{R}$ and $\mathcal{W}$ respectively.
2) Obtaining investments. We then solve (II-B) for the given representative days to obtain investment decisions, ${x}$, and the corresponding investment cost, ${c}^{\top}{x}$.
3) Evaluating operational costs. Given the investment decisions $x$ from the second step, we solve the resulting recourse problem (II-A) for the full planning horizon, $\mathcal{T}$, for all scenarios in $\Omega^{\text{val}}$. This yields a set of operational decisions and corresponding operational costs for all validation scenarios. Averaging these values gives $f(\theta)={c}^{\top}{x}+\frac{1}{|\Omega^{\text{val}}|}\sum_{\omega\in\Omega^{\text{val}}}\sum_{t\in\mathcal{T}}d_{t}^{\top}y_{t}^{\omega}$, an upper bound (i.e., feasible solution) for (1a) instantiated on $\Omega^{\text{val}}$.
Regarding computational runtime, solutions for $\text{RPC}(\theta)$ can be obtained quickly with approximate algorithms such as PAM [30]. The computational burden in Step (3) is also likely to be low as it requires solving $|\Omega^{\text{val}}|$ recourse problems, which are linear programs in our case. Additionally, Step (3) can easily be parallelized to accelerate the process. In most cases, Step (2) of the function evaluation will require the greatest runtime as it involves solving a mixed-integer linear program (II-B) with $\mathcal{O}(|\mathcal{T}_{R}|)$ second-stage variables and constraints. Consequently, the worst-case runtime of one function evaluation depends on the maximum value that $\theta_{R}$ (i.e., the number of representative days) takes in $\Theta$.
III-B Selecting Candidate Hyperparameters
@@
1) Instantiating a prior. We instantiate a prior predictive covariance function, or kernel function, $\Sigma\mathrel{\mathop{:}}\Theta\times\Theta\to\mathbb{R}$, which yields the GP prior distribution
$f^{1\mathrel{\mathop{:}}j}\sim\mathcal{N}(0,\Sigma({\theta}^{1\mathrel{\mathop{:}}j},{\theta}^{1\mathrel{\mathop{:}}j})),$ (5)
where ${\theta}^{1\mathrel{\mathop{:}}j}$ denotes $\theta^{1},\dots,\theta^{j}$ and the kernel function is defined as:
@@
where $\Gamma(\cdot)$ and $K_{5/2}(\cdot)$ are the gamma function and modified Bessel function respectively [33]. This is a common choice of kernel function [33, 32] and is the default setting in BoTorch [34], the package that we used in our experiments. Note that in (5) we assume $f^{1\mathrel{\mathop{:}}j}$ to have zero mean, which we enforce in practice through standardization.
@@
$f\mid\theta,\theta^{1\mathrel{\mathop{:}}j},f^{1\mathrel{\mathop{:}}j}\sim\mathcal{N}(\mu(\theta),\sigma^{2}(\theta)),$
@@
$\mu(\theta)$ $=\Sigma(\theta,\theta^{1\mathrel{\mathop{:}}j})\Sigma(\theta^{1\mathrel{\mathop{:}}j},\theta^{1\mathrel{\mathop{:}}j})^{-1}f^{1\mathrel{\mathop{:}}j}$ (6)
$\sigma^{2}(\theta)$ $=\Sigma(\theta,\theta)-\Sigma(\theta,\theta^{1\mathrel{\mathop{:}}j})\Sigma(\theta^{1\mathrel{\mathop{:}}j},\theta^{1\mathrel{\mathop{:}}j})^{-1}\Sigma(\theta^{1\mathrel{\mathop{:}}j},\theta)$ (7)
@@ IV Computational Study
Using BoTorch [34], we implement two BO approachs with different GP hyperparameters, $\beta=10$ and $\beta=2$ (Eq. 8). We denote these two methods by $\text{BO}_{10}$ and $\text{BO}_{2}$ respectively. We conduct four trials, each of which corresponds to an initialization with $N_{0}=20$ random function evaluations followed by 80 iterations of our BO-assisted approach. The resulting hyperparameters are evaluated according to their corresponding validation cost, $f^{\text{val}}$ (Sec. III-A), and test cost, $f^{\text{test}}$, defined to be the average cost incurred over the set of test projections. We also compare our approach to a random search heuristic that evaluates hyperparameters $\theta$ sampled uniformly at random from $\Theta$.
@@
TABLE I: Hyperparameters, objective values, and improvement over baseline costs for each method (corresponding to lowest test cost across four trials). Percentage improvement is calculated with respect to the best performing baseline solution for the respective scenario set (i.e., Base. 1 for the validation set and Base. 2 for the test set). Dashes denote zero values.
@@
(b) Top: histograms of daily parameter values observed in the sample set (blue) and in the rep. day sets corresponding to Table I (shown as kernel density estimates). Values are averaged over hours and nodes for all parameters except gas load. Bottom: quantile-quantile plots comparing the empirical distribution of the sample set to the identified rep. day sets.
@@ IV-A Cost Comparison
Table I shows the best hyperparameter configuration identified for each method over all trials and its corresponding cost and improvement over the baselines. Interestingly, test costs are lower than validation costs across most experiments. This is a reflection of the relative harder-to-meet supply-demand projections that exist in the validation set rather than a statistical bias intrinsic to the optimization procedure. Importantly, all search methods improve over the baseline test costs by at least 3% despite using fewer representative days. This demonstrates the necessity of tuning RPC hyperparameters over instantiating a large surrogate problem using generic (i.e., untuned) distance metrics.
IV-B Hyperparameter Values
@@ IV-C New Assets and Utilization
TABLE II: Installed VRE capacity (GW), VRE utilization rate (%), and storage investment (\$ billion). Utilization rate is calculated as average hourly generation divided by installed capacity.
@@
Compared to Base. 1 and Base. 2, all search methods invest less in expanding generation capacity but more in energy storage infrastructure (Tab. II). Nevertheless, their resulting investments can substantially reduce costs from load shedding in the power system when compared to Base. 1, and in the case of $\text{BO}_{2}$, operate with zero load shedding (similar to Base. 2) as a result of including two extreme days (Fig. 8). Interestingly, Tab. II shows that the utilization rate of VRE assets is generally higher for the search approaches than the baselines. In particular, $\text{BO}_{10}$, which weighted solar CFs more heavily in clustering, is able to dispatch more power from the same installed solar capacity through optimizing geographical distribution of plant locations. These findings reflect the cost savings obtained by our BO-assisted approach resulting from higher investment in storage coupled with establishment of relatively fewer VRE assets that can nevertheless be dispatched more effectively through strategic positioning.
V Concluding Remarks

### v4

@@
Learning-assisted Stochastic Capacity Expansion Planning: A Bayesian Optimization Approach
Aron Brenner Rahman Khorramfar Dharik Mallapragada Saurabh Amin
(July 9, 2024)
@@
1 Introduction
1.1 Motivation
@@
1.2 Time Series Aggregation
@@
Figure 1: Effect of varying the number of representative days on the CEP objective evaluated over 14 out-of-sample supply-demand projections using a single projection solution for a joint power-gas CEP (Sec. 4).
1.3 Contributions
We consider two-stage stochastic CEP for coordinated planning of multiple energy vectors. We develop a learning-assisted approach to compute investment decisions that yield the lowest combined investment and operational cost in expectation over multiple year-long projections of energy demand and VRE availability (Sec. 2). Rather than solve an extensive form stochastic program, which is intractable due to the large-scale and combinatorial nature of the problem, our approach designs and solves a surrogate CEP instantiated using a smaller set of representative periods. Note, however, that there can be many potential surrogate CEPs, each yielding different planning outcomes. Our key contribution is to leverage a Bayesian optimization (BO) approach to efficiently search the space of RPC hyperparameters. We demonstrate that these automatically tuned hyperparameters help identify planning outcomes with the lowest expected total cost (investment plus operational) over out-of-sample supply-demand projections (Sec. 3).
@@ 1.3 Contributions
Our BO-assisted approach can be viewed as a “learning to configure” approach [22] in that it learns to select hyperparameters specifying the time series aggregation heuristic for surrogate model construction. In this regard, our approach is distinct from recent works that use offline learning to improve optimization heuristics [23, 24] as it can be deployed without pre-training on a large number of problem instances (Fig. 3). Specifically, our approach quickly identifies low-cost investment decisions by learning to identify promising hyperparameter configurations for a single problem instance through an iterative process of instantiating surrogate problems, solving them, and evaluating the resulting investment decisions over a range of validation supply-demand projections. We demonstrate the effectiveness of this approach for joint power-gas generation and transmission expansion planning for New England with 20 yearly supply-demand projections (Sec. 4).
2 Problem Formulation
2.1 Capacity Expansion Problem
@@
We assume that (2) is feasible for any first-stage decisions, $x$, i.e., load shedding is unconstrained and incurred as a cost. Additionally, we note that each period, $t\in\mathcal{T}$, does not necessarily correspond to the planning resolution of energy systems. For instance, while hourly time resolution is usually used for electric power planning, daily operating resolution can be more suitable for gas system planning [3]. Following [13, 26, 27], we assume representative periods to denote daily time spans.
2.2 Surrogate Capacity Expansion Problem
@@
2.3 Representative Period Clustering Problem
The choice of representative days is crucial to achieving a computationally light surrogate model that well-approximates (1). In other words, the investment decisions $x$ resulting from $\text{CEP}_{\text{S}}(\mathcal{T}_{R},\mathcal{W})$ should approximately minimize (1a). Existing literature, including [29, 26, 14, 12], propose clustering-based approaches for identifying such representative periods. Here, we consider the following k-medoids clustering model, which we refer to as the representative period clustering (RPC) problem:
@@ 2.3 Representative Period Clustering Problem
where $\mathcal{C}\coloneqq(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ denotes the representative day clusters with corresponding medoids $(m_{1},\dots,m_{k})$, $D(m_{i},t)$ denotes the “distance” between medoid $m_{i}$ and period $t$, and $\mathcal{T}_{E}\subset\mathcal{T}_{R}$ denotes a set of “extreme days,” or days exhibiting particularly hard-to-satisfy demands, with corresponding weights $w_{i}=1$ for all $i\in\mathcal{T}_{E}$. One can apply an out-of-the-box clustering algorithm [30] to solve (4) and obtain the representative day set and weights respectively as $\mathcal{T}_{R}=(m_{1},\dots,m_{\theta_{R}})$ and $\mathcal{W}=(\frac{1}{|\mathcal{S}|}|\mathcal{C}_{1}|,\dots,\frac{1}{|\mathcal{S}|}|\mathcal{C}_{\theta_{R}}|)$. Here, we focus on the central issue of selecting hyperparameters for the RPC problem.
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We consider $\theta_{B}\in[0,1]$ to be a hyperparameter that weighs the relative importance of demand (contained in ${q}_{t}^{\omega}$) and supply-side features such as renewable capacity factors (contained in ${B}_{t}^{\omega}$). Importantly, different choices of $\theta_{B}$ impact RPC outcomes by influencing the daily supply- or demand- patterns captured in the representative day set. Selecting $\theta_{B}$ is not obvious. More generally, more than two groups of parameters may need to be weighed when multiple VRE technologies and energy vectors are considered, which is the case for our computational study (Sec. 4).
@@
Figure 5: Pearson correlations between supply-demand parameter groups for our computational study (Sec. 4). Each row/column of pixels corresponds to one hour of the day (averaged over system nodes) except for NG, where pixels correspond to daily nodal loads. For each pair of parameter groups, pixel $(i,j)$ shows the correlation of parameter $i$ in the first group with parameter $j$ in the second group.
@@
3 Solution Approach
To address this challenge systematically and for a range of planning contexts, we propose searching over a continuous space of RPC hyperparameters using Bayesian optimization (BO), a derivative-free search strategy for black-box functions that has been applied with success in hyperparameter tuning for ML applications [32]. At a high level, our BO-assisted approach learns to minimize $f(\cdot)$, the function mapping RPC hyperparameters to CEP objective values, by performing a series of function evaluations. As is the case in many BO applications, $f(\cdot)$ does not admit an analytic expression and can only be evaluated through a series of complex operations (i.e., solving optimization problems). In each iteration of our approach, a function estimate – a Gaussian process (GP) regression model – is re-estimated using the most recent function evaluation, and new candidate hyperparameters are identified according to their estimated potential for yielding a low CEP objective and improving the GP model estimate. Fig. 7 illustrates our solution approach and shows how hyperparameters $\theta$ are evaluated (Sec. 3.1) and how subsequent candidate hyperparameters are selected (Sec. 3.2) in each iteration.
3.1 Function Evaluation
@@
1) Clustering the sample set. We solve (4) using a sample set constructed from $\Omega^{\text{train}}$ to identify a set of representative days (i.e., cluster medoids) and their corresponding weights, $\mathcal{T}_{R}$ and $\mathcal{W}$ respectively.
2) Obtaining investments. We then solve (3) for the given representative days to obtain investment decisions, ${x}$, and the corresponding investment cost, ${c}^{\top}{x}$.
3) Evaluating operational costs. Given the investment decisions $x$ from the second step, we solve the resulting recourse problem (2) for the full planning horizon, $\mathcal{T}$, for all scenarios in $\Omega^{\text{val}}$. This yields a set of operational decisions and corresponding operational costs for all validation scenarios. Averaging these values gives $f(\theta)={c}^{\top}{x}+\frac{1}{|\Omega^{\text{val}}|}\sum_{\omega\in\Omega^{\text{val}}}\sum_{t\in\mathcal{T}}d_{t}^{\top}y_{t}^{\omega}$, an upper bound (i.e., feasible solution) for (1a) instantiated on $\Omega^{\text{val}}$.
Regarding computational runtime, solutions for $\text{RPC}(\theta)$ can be obtained quickly with approximate algorithms such as PAM [30]. The computational burden in Step (3) is also likely to be low as it requires solving $|\Omega^{\text{val}}|$ recourse problems, which are linear programs in our case. Additionally, Step (3) can easily be parallelized to accelerate the process. In most cases, Step (2) of the function evaluation will require the greatest runtime as it involves solving a mixed-integer linear program (3) with $\mathcal{O}(|\mathcal{T}_{R}|)$ second-stage variables and constraints. Consequently, the worst-case runtime of one function evaluation depends on the maximum value that $\theta_{R}$ (i.e., the number of representative days) takes in $\Theta$.
3.2 Selecting Candidate Hyperparameters
@@
1) Instantiating a prior. We instantiate a prior predictive covariance function, or kernel function, $\Sigma:\Theta\times\Theta\to\mathbb{R}$, which yields the GP prior distribution
$f^{1:j}\sim\mathcal{N}(0,\Sigma({\theta}^{1:j},{\theta}^{1:j})),$ (5)
where ${\theta}^{1:j}$ denotes $\theta^{1},\dots,\theta^{j}$ and the kernel function is defined as:
@@
where $\Gamma(\cdot)$ and $K_{5/2}(\cdot)$ are the gamma function and modified Bessel function respectively [33]. This is a common choice of kernel function [33, 32] and is the default setting in BoTorch [34], the package that we used in our experiments. Note that in (5) we assume $f^{1:j}$ to have zero mean, which we enforce in practice through standardization.
@@
$f\mid\theta,\theta^{1:j},f^{1:j}\sim\mathcal{N}(\mu(\theta),\sigma^{2}(\theta)),$
@@
$\mu(\theta)$ $=\Sigma(\theta,\theta^{1:j})\Sigma(\theta^{1:j},\theta^{1:j})^{-1}f^{1:j}$ (6)
$\sigma^{2}(\theta)$ $=\Sigma(\theta,\theta)-\Sigma(\theta,\theta^{1:j})\Sigma(\theta^{1:j},\theta^{1:j})^{-1}\Sigma(\theta^{1:j},\theta)$ (7)
@@ 4 Computational Study
Using BoTorch [34], we implement two BO approachs with different GP hyperparameters, $\beta=10$ and $\beta=2$ (Eq. 8). We denote these two methods by $\text{BO}_{10}$ and $\text{BO}_{2}$ respectively. We conduct four trials, each of which corresponds to an initialization with $N_{0}=20$ random function evaluations followed by 80 iterations of our BO-assisted approach. The resulting hyperparameters are evaluated according to their corresponding validation cost, $f^{\text{val}}$ (Sec. 3.1), and test cost, $f^{\text{test}}$, defined to be the average cost incurred over the set of test projections. We also compare our approach to a random search heuristic that evaluates hyperparameters $\theta$ sampled uniformly at random from $\Theta$.
@@
Table 1: Hyperparameters, objective values, and improvement over baseline costs for each method (corresponding to lowest test cost across four trials). Percentage improvement is calculated with respect to the best performing baseline solution for the respective scenario set (i.e., Base. 1 for the validation set and Base. 2 for the test set). Dashes denote zero values.
@@
(b) Top: histograms of daily parameter values observed in the sample set (blue) and in the rep. day sets corresponding to Table 1 (shown as kernel density estimates). Values are averaged over hours and nodes for all parameters except gas load. Bottom: quantile-quantile plots comparing the empirical distribution of the sample set to the identified rep. day sets.
@@ 4.1 Cost Comparison
Table 1 shows the best hyperparameter configuration identified for each method over all trials and its corresponding cost and improvement over the baselines. Interestingly, test costs are lower than validation costs across most experiments. This is a reflection of the relative harder-to-meet supply-demand projections that exist in the validation set rather than a statistical bias intrinsic to the optimization procedure. Importantly, all search methods improve over the baseline test costs by at least 3% despite using fewer representative days. This demonstrates the necessity of tuning RPC hyperparameters over instantiating a large surrogate problem using generic (i.e., untuned) distance metrics.
4.2 Hyperparameter Values
@@ 4.3 New Assets and Utilization
Table 2: Installed VRE capacity (GW), VRE utilization rate (%), and storage investment (\$ billion). Utilization rate is calculated as average hourly generation divided by installed capacity.
@@
Compared to Base. 1 and Base. 2, all search methods invest less in expanding generation capacity but more in energy storage infrastructure (Tab. 2). Nevertheless, their resulting investments can substantially reduce costs from load shedding in the power system when compared to Base. 1, and in the case of $\text{BO}_{2}$, operate with zero load shedding (similar to Base. 2) as a result of including two extreme days (Fig. 9). Interestingly, Tab. 2 shows that the utilization rate of VRE assets is generally higher for the search approaches than the baselines. In particular, $\text{BO}_{10}$, which weighted solar CFs more heavily in clustering, is able to dispatch more power from the same installed solar capacity through optimizing geographical distribution of plant locations. These findings reflect the cost savings obtained by our BO-assisted approach resulting from higher investment in storage coupled with establishment of relatively fewer VRE assets that can nevertheless be dispatched more effectively through strategic positioning.
5 Concluding Remarks